OpenStudy (anonymous):

anti-derivative of x^3sqrt(x^2-1) using a u substitution

ganeshie8 (ganeshie8):

first thing you should try is to sub the stuff inside radical

ganeshie8 (ganeshie8):

sub \(\large u = x^2-1\)

OpenStudy (anonymous):

did that

OpenStudy (anonymous):

take du=2x dx, dx= 1/2x du

ganeshie8 (ganeshie8):

yes, we can do better : \(\large u = x^2 - 1 \implies x^2 = u-1\) \(\large du = 2x dx\) \(\large xdx = \frac{1}{2}du\)

ganeshie8 (ganeshie8):

\(\large \int x^3\sqrt{x^2-1}~ dx\) \(\large \int x^2. x \sqrt{x^2-1}~ dx\) \(\large \int x^2 \sqrt{x^2-1} ~xdx\)

ganeshie8 (ganeshie8):

substitute now ^

ganeshie8 (ganeshie8):

the integral becomes : \(\large \int (u+1)\sqrt{u} ~ \dfrac{du}{2}\)

ganeshie8 (ganeshie8):

should be trivial to evaluate ?

OpenStudy (anonymous):

should that not be du/2x

ganeshie8 (ganeshie8):

that x got stuck with the "dx" on the left hand side itself

OpenStudy (anonymous):

ok so, now what? this is exactly were i got stuck on my own lol. do i sub back for u to get an equation in the form of x^2sqrt(x^2-1)du/2?

ganeshie8 (ganeshie8):

\(\large \int (u+1)\sqrt{u} ~ \dfrac{du}{2}\)

ganeshie8 (ganeshie8):

you're fine till that step ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

multiply

ganeshie8 (ganeshie8):

\(\large \dfrac{1}{2}\int (u\sqrt{u}+\sqrt{u} )~ du\)

ganeshie8 (ganeshie8):

\(\large \dfrac{1}{2}\int (u^{\frac{3}{2}}+u^{\frac{1}{2}})~ du\)

ganeshie8 (ganeshie8):

use below formula to evaluate : \[\large \int x^n dx = \dfrac{x^{n+1}}{n+1}+C\]

OpenStudy (anonymous):

is that a calc 2 formula? this is calc 1 and i dont believe we were given that to evaluate our types of intergrals

ganeshie8 (ganeshie8):

thats calc1 formula

OpenStudy (anonymous):

ok ok thats familiar

ganeshie8 (ganeshie8):

good :)

OpenStudy (anonymous):

by doing that i get 1/2[(x^2-1)^5/2)/5/2+(x^2-1)^3/2)/3/2))

ganeshie8 (ganeshie8):

looks good ! don't forget the constant C

OpenStudy (anonymous):

the correct answer according to wolfram was supposed to be \[[(x ^{2}-1)^3/2 *(3x^2 +2)]/15 \]

OpenStudy (anonymous):

thanks so much. ill be posting another similar substitution in a few minutes once i jot this down in my note book

ganeshie8 (ganeshie8):

np :)